# Topological photon pumping in quantum optical systems

### v2

@@ 3 Long-range hopping in quantum optics: the extended Rice-Mele model
In this paper we focus on the realization of topological pumping using quantum optics platforms and, in this Section, we present the general equation that describes the dynamics of these systems. We then introduce a mapping onto a generalized version of the Rice-Mele model with long-range hopping rates, which we refer to as the extended Rice-Mele model, and discuss its topological properties.
All three quantum optics platforms we discuss here share the following description (see Fig. 1a): An ensemble of $N$ two-level emitters in the Lamb-Dicke regime with ground state $\left|g\right>$ and excited state $\left|e\right>$ separated by an energy $\hbar\omega_{0}$, are coupled to a common environment  [^1] Here, the zero-temperature radiation field, either in free-space or with imposed boundary conditions that arise from the presence of a nanophotonic structure, such as a nanofiber. Within the Born-Markov and secular approximations, the quantum master equation that describes the evolution of the emitter’s internal degrees of freedom is
@@
with the spin-$1/2$ ladder operators $\hat{\sigma}_{i}=\left|g_{i}\right>\left<e_{i}\right|$ and $\hat{\sigma}_{i}=\left|e_{i}\right>\left<g_{i}\right|$. This Hamiltonian drives an exchange (dipole-dipole) interaction between the emitters, which occurs at a rate $V_{ij}$. The strength and range of these exchange rates will be determined by the system at hand. The second term, $\hat{H}_{\Delta}$, is an on-site potential which leads to an offset $2\hbar\Delta$ between even and odd sites (see Fig. 1b). The dissipation in the system (second term of the master equation) also depends on the system at hand, and its action on the dynamics will be considered separately for each system in Sec. 5.1.
@@
In the next step, we make the bipartite lattice explicit by denoting $\hat{a}_{q}=\hat{f}_{2q-1}$ and $\hat{b}_{q}=\hat{f}_{2q}$ where $q=1\dots N/2$ labels the unit cells. Accordingly, we categorize $V_{ij}$ into three different hopping rates: from sublattice $A$ to $B$, from $B$ to $A$, and within the same sublattice, which we denote $J^{\prime}_{2p-1}$, $J_{2p-1}$ and $J_{2p}$, respectively, where $p$ denotes again the unit cell (see Fig. 4a). The extended Rice-Mele Hamiltonian is then given by
@@ 3.1 Topological properties of the extended Rice-Mele model
that contain the intra-sublattice and inter-sublattice hopping rates respectively. Diagonalizing $h(k)$ we obtain the dispersion relation $E_{\pm}(k)=n_{0}(k)\pm\sqrt{|n(k)|^{2}+\Delta^{2}}$ for the upper ($+$) and lower (--) band, respectively. The extended Rice-Mele Hamiltonian with $\Delta\neq 0$ does not host topological phases. However, for vanishing energy offset $\Delta=0$, the Hamiltonian reduces to the extended SSH Hamiltonian [38]. This Hamiltonian can be allocated to the BDI symmetry class, which possesses a $\mathbb{Z}$-type topological invariant in one dimension, like the standard SSH model [23], when the sublattice symmetry is conserved, i.e. when $n_{0}(k)=0$, which can be accomplished by setting $J_{2p}=0$ for all $p$.
@@
for $k=0$, making the argument of $n(k)$ undefined and the energy gap zero. Note, that there are further specific fine-tuned combinations of the parameters $J_{2p-1}$ and $J^{\prime}_{2p-1}$ that lead to topological degeneracy points at other points of the FBZ. However, in the pumping cycles and physical systems that we propose in this paper the gap only closes at $k=\pm\pi/a$. Thus, we introduce the extended hopping rates
@@ 4 Topological photon pumping in quantum optical systems
Having set up the stage for the realization of dispersionless topological pumping in the extended Rice-Mele model (see Fig. 4a), we now examine three quantum optics platforms of current experimental relevance. These platforms enable the pumping of photons stored in the excited state of two-level emitters and encompass: a Rydberg lattice, a dense lattice gas of atoms in low-lying electronic states, and atoms coupled to a nanophotonic waveguide as sketeched in Fig. 4b - d. These three platforms all present long-range excitation hopping rates, but the different shape and strength of the interactions lead to differences in the effective range of the rates. E.g. in a Rydberg lattice system the nearest neighbor hopping rates dominate the dynamics, while in the waveguide system all atoms are coupled to all other atoms with similar strength, such that excitation hopping stretches across all distances.
Figure 4: Realizations of the extended Rice-Mele model. a: In the extended Rice-Mele model, we consider in general non-zero hopping rates between all sites. Here, an example for a system with $N=8$ sites. b: In the Rydberg system, the two sublattices are separated by a perpendicular distance $h$ and a horizontal offset $b(t)$, which can be changed in time to realize the photon pumping. Here, the nearest neighbor hoppings dominate and we exploit the dependence of the interactions on dipole moment orientation (right panel) to make the intra-sublattice hoppings ($J_{2p}$) negligible. c: Similarly, on a dense chain of atoms excited to low-lying electronic states, $J_{2}$ is put to zero by fixing a value of $a$ that depends on the angle $\theta_{\mathrm{d}}$ of the dipoles, while the rest of the hopping rates (determined by $V_{ij}$) are non-zero in general. d: In a lattice of atoms coupled to a waveguide (here a cylindrical nanofiber), the atoms of sublattice $A$ are fixed forming an helix around the fiber, while the atoms of sublattice $B$ change their relative angle to sublattice $A$, $\phi^{\prime}_{1}(t)$, time-dependently to realize the pumping. Sublattice symmetry is preverved by choosing $a=\pi/\beta$, where $\beta$ is the wave vector of the light propagating inside the fiber.
@@ 4.1 Short-range hopping: Rydberg atoms
Figure 5: Photon pumping on a Rydberg chain. a: Dependence of the extended hopping rates $\bar{J}^{\prime}$ and $\bar{J}$ with the ratio $b/a$ (all rates are normalized to the maximal rate $\bar{J}_{\text{max}}\equiv|\text{max}\{\bar{J}(t)\}|$ during the cycle). The gray shaded area represents $\delta\bar{J}<0$, where the winding number of the extended SSH model is $\nu=1$ (TP: topological phase). b: Variation of the intracell distance $b(t)$ during the pumping cycle. Importantly, twice during the cycle the extended hopping rates coincide, such that $\delta\bar{J}=0$. c: We vary the energy offset as $\Delta(t)=\Delta_{\mathrm{max}}\sin(2\pi t/T+\pi/2)$ such that the path in parameter space encircles the topological degeneracy point ensuring the displacement of the center of mass after a completed cycle. d: The left column shows the time-integrated Berry curvature of the lower band ${\cal W}_{k}^{-}$ and quasi-momentum distribution density $|f(k)|^{2}$ (in arbitrary units). The right column shows the spatial probability density of the wave packet in real space (on the $B$ sublattice), initially (dashed line) and after its numerical evolution for ten pumping cycles (solid line). The width of the Gaussian wave packet in momentum space, $w_{k}$, is doubled from the upper ($w_{k}a=2\pi/100$) to the lower row ($w_{k}a=2\pi/50$), leaving the displacement of the photon’s center of mass $\Delta x=10\cdot 0.966a$ unchanged, but leading to an increase of the dispersion, such that the fidelity per cycle ${\cal F}(T)=0.999$ and after ten cycles ${\cal F}(10T)=0.988$ decreases to 0.997 and 0.964, respectively.
The first system we consider is an ensemble of Rydberg atoms trapped in two lattices with lattice constant $a$ and offset by a distance $b$ with respect to each other as shown in Fig. 4[42]. Atoms excited to a Rydberg state are known for possessing exaggerated properties, such as long lifetimes and, most importantly, extremely large transition dipole moments between neighboring Rydberg states which can lead to strong long-range dipole-dipole interactions [69, 70]. The dynamics of the system is determined by the Hamiltonian (8), where both ground and excited states of the two-level system are Rydberg states. The dipole-dipole interaction is here given generally by the expression
@@
where $\mathbf{r}_{ij}=r_{ij}\hat{\mathbf{r}}_{ij}$ is the distance between the atoms, $\mathbf{d}=d\hat{\mathbf{d}}$ is the dipole moment of the transition, and $\cos{\theta_{ij}}=\hat{\mathbf{r}}_{ij}\cdot\hat{\mathbf{d}}$. To avoid breaking sublattice symmetry, we set the intra-sublattice rates $J_{2p}$ to zero by aligning the dipoles such that they form an angle $\theta_{m}=\arccos(1/\sqrt{3})$ with the unit separation vector between atoms of the same sublattice (see Fig. 4b). Moreover, the dipole-dipole interactions (excitation hopping rates) between neighboring atoms (lattice sites) dominate the dynamics, such that $\bar{J}^{\prime}\approx J_{1}^{\prime}$ and $\bar{J}\approx J_{1}$ with
@@
Here, $r_{1}^{\prime}=\sqrt{b^{2}+h^{2}}$, $r_{1}=\sqrt{(a-b)^{2}+h^{2}}$, $\cos\theta_{1}^{\prime}=\left(b\cos\theta_{m}-h\sin\theta_{m}\right)/r_{1}^{\prime}$ and $\cos\theta_{1}=\left[(a-b)\cos\theta_{m}+h\sin\theta_{m}\right]/r_{1}$. Adding an on-site potential offset between the even and odd sites, the Hamiltonian is very close to the standard Rice-Mele model (4). Note, however, that this is only an approximation, and the the actual long-ranged character of the interactions in the Rydberg system is fully taken into account in the following in our numerical calculations.
@@ 4.1.1 Photon pumping on a Rydberg chain
Figure 6: Photon pumping on a two-level atom chain. a: Dependence of the extended hopping rates $\bar{J}^{\prime}$ and $\bar{J}$ with the ratio $b/a$ (normalized to $\bar{J}_{\text{max}}\equiv|\text{max}\{\bar{J}(t)\}|$ during the cycle). The gray shaded area represents $\delta\bar{J}<0$, where the winding number of the extended SSH model is $\nu=1$ (TP: topological phase). b: Variation of the intracell distance $b(t)$ during the pumping cycle. c: We vary the energy offset as $\Delta(t)=\Delta_{\mathrm{max}}\sin(2\pi t/T+\pi/2)$ such that the path in parameter space encircles the topological degeneracy point ensuring the displacement of the center of mass after a completed cycle. d: The left column shows the time-integrated Berry curvature of the lower band ${\cal W}_{k}^{-}$ and quasi-momentum distribution density $|f(k)|^{2}$ (in arbitrary units). The right column shows the spatial probability density of the wave packet in real space (on the $B$ sublattice), initially (dashed line) and after its numerical evolution for ten pumping cycles (solid line). The width of the Gaussian wave packet in momentum space $w_{k}$ (the same as in Fig. 5) is doubled from the upper to the lower row, leaving the displacement of the photon’s center of mass $\Delta x=10\cdot 0.803a$ unchanged, but leading to an increase of the dispersion, such that the fidelity per cycle ${\cal F}(T)=0.999$ and after ten cycles ${\cal F}(10T)=0.986$ decreases to 0.986 and 0.890, respectively.
We simulate numerically the dynamics by solving the Schrödinger equation with the time-dependent Hamiltonian $\hat{H}(t)$ (LABEL:ExtendedRM). We choose an initial wave packet on the lower band, with a Gaussian quasi-momentum distribution $|f(k)|^{2}$ of width $w_{k}$. Here, we are able to find an optimized path in parameter space that leads to a particularly flat time-integrated Berry curvature across all the FBZ. As one can also see in this figure, the cyclic pumping indeed leaves the shape of the photon mostly unchanged except for its average position, which is displaced by an amount $\Delta x$. This displacement is in agreement with the one based on the time-integrated Berry curvature and the quasi-momentum distribution predicted by Equation (2). Notably, here we are able to achieve a fidelity per cycle well above $99.9\%$, which goes down as we increase the width of the wave packet in $k$-space and dispersion plays a more prominent role.
@@ 4.2 Long-range hopping: Atoms in free space
where $\gamma=\frac{d^{2}k_{0}^{3}}{3\pi\epsilon_{0}\hbar}$ is the single atom spontaneous decay rate, and $k_{0}=\omega_{0}/c=2\pi/\lambda_{0}$ is the wave number of the transition, respectively. This interaction can no longer be realistically approximated to be nearest neighbor only, as it was the case in the Rydberg system. Thus, the corresponding extended Rice-Mele model (LABEL:ExtendedRM) contains all hopping rates $J^{\prime}_{2p-1}$, $J_{2p-1}$ and $J_{2p}$. While the sublattice symmetry is in general broken in this system, we minimize this effect by choosing the distance $a$ between neighboring atoms of the same sublattice at a value $a(\theta_{d})$, where $\theta_{d}$ is the angle between the atomic dipoles and the chain, i.e. $\cos{\theta_{d}}=\hat{\mathbf{r}}_{ij}\cdot\hat{\mathbf{d}}$, such that $J_{2}=0$ (see Fig. 4c for the $\theta_{d}=\pi/2$ case). Given that the hopping rates slowly decrease with the distance between atoms, the values of the subsequent symmetry-breaking hopping rates $J_{4}$, $J_{6}$, are almost negligible, and the sublattice symmetry is only slightly broken.

### v3

@@ 3 Long-range hopping in quantum optics: the extended Rice-Mele model
In this paper we focus on the realization of topological pumping using quantum optics platforms and, in this Section, we present the general equation that describes the dynamics of these systems. We then introduce a mapping onto a generalized version of the Rice-Mele model with long-range hopping parameters, which we refer to as the extended Rice-Mele model, and discuss its topological properties.
All three quantum optics platforms we discuss here share the following description (see Fig. 1a): An ensemble of $N$ two-level emitters in the Lamb-Dicke regime with ground state $\left|g\right>$ and excited state $\left|e\right>$ separated by an energy $\hbar\omega_{0}$, are coupled to a common environment [^1] Here, the zero-temperature radiation field, either in free-space or with imposed boundary conditions that arise from the presence of a nanophotonic structure, such as a nanofiber.. Within the Born-Markov and secular approximations, the quantum master equation that describes the evolution of the emitter’s internal degrees of freedom is
@@
with the spin-$1/2$ ladder operators $\hat{\sigma}_{i}=\left|g_{i}\right>\left<e_{i}\right|$ and $\hat{\sigma}_{i}=\left|e_{i}\right>\left<g_{i}\right|$. This Hamiltonian drives an exchange (dipole-dipole) interaction between the emitters. The strength and range of the exchange, encoded in the matrix elements $V_{ij}$, will be determined by the system at hand. The second term, $\hat{H}_{\Delta}$, is an on-site potential which leads to an offset $2\hbar\Delta$ between even and odd sites (see Fig. 1b). The dissipation in the system (second term of the master equation) also depends on the system at hand, and its action on the dynamics will be considered separately for each system in Sec. 5.1.
@@
In the next step, we make the bipartite lattice explicit by denoting $\hat{a}_{q}=\hat{f}_{2q-1}$ and $\hat{b}_{q}=\hat{f}_{2q}$ where $q=1\dots N/2$ labels the unit cells. Accordingly, we categorize $V_{ij}$ into three different hopping parameters: from sublattice $A$ to $B$, from $B$ to $A$, and within the same sublattice, which we denote $J^{\prime}_{2p-1}$, $J_{2p-1}$ and $J_{2p}$, respectively, where $p$ denotes again the unit cell (see Fig. 4a). The extended Rice-Mele Hamiltonian is then given by
@@ 3.1 Topological properties of the extended Rice-Mele model
that contain the intra-sublattice and inter-sublattice hopping parameters respectively. Diagonalizing $h(k)$ we obtain the dispersion relation $E_{\pm}(k)=n_{0}(k)\pm\sqrt{|n(k)|^{2}+\Delta^{2}}$ for the upper ($+$) and lower (--) band, respectively. The extended Rice-Mele Hamiltonian with $\Delta\neq 0$ does not host topological phases. However, for vanishing energy offset $\Delta=0$, the Hamiltonian reduces to the extended SSH Hamiltonian [34, 38]. This Hamiltonian can be allocated to the BDI symmetry class, which possesses a $\mathbb{Z}$-type topological invariant in one dimension, like the standard SSH model [23], when the sublattice symmetry is conserved, i.e. when $n_{0}(k)=0$, which can be accomplished by setting $J_{2p}=0$ for all $p$. Note, however, that even though we impose here, where possible, sublattice symmetry, the model described by (LABEL:ExtendedRM) is inversion symmetric and hence $J_{2p}=0$ is not a necessary condition for the pumping to take place.
@@
for $k=0$, making the argument of $n(k)$ undefined and the energy gap zero. Note, that there are further specific fine-tuned combinations of the parameters $J_{2p-1}$ and $J^{\prime}_{2p-1}$ that lead to topological degeneracy points at other points of the FBZ. However, in the pumping cycles and physical systems that we propose in this paper the gap only closes at $k=\pm\pi/a$. Thus, we introduce the extended hopping parameters
@@ 4 Topological photon pumping in quantum optical systems
Having set up the stage for the realization of dispersionless topological pumping in the extended Rice-Mele model (see Fig. 4a), we now examine three quantum optics platforms of current experimental relevance. These platforms enable the pumping of photons stored in the excited state of two-level emitters and encompass: a Rydberg lattice, a dense lattice gas of atoms in low-lying electronic states, and atoms coupled to a nanophotonic waveguide as sketeched in Fig. 4b - d. These three platforms all present long-range interactions, but the different shape and strength of the couplings lead to differences in their effective range. E.g. in a Rydberg lattice system the nearest neighbor hopping parameters dominate the dynamics, while in the waveguide system all atoms are coupled to all other atoms with similar strength, such that excitation hopping stretches across all distances.
In all cases, we numerically simulate the dynamics of the Schrödinger equation with a time-dependent Hamiltonian of the form (LABEL:ExtendedRM) for a system with open boundary conditions. Note that the cycle period $T$ needs to be sufficiently long to ensure adiabaticity. This is fulfilled when $T\gg\hbar/\Delta E$, where $\Delta E$ is the minimum energy gap between neighboring bands during the cycle. Conversely, our transport protocols aim to minimize dispersion, requiring that the cycle is fast enough to prevent non-linearities in the Berry phase $\gamma(k)$ from introducing wave packet dispersion Throughout the paper, we have found cycle periods that approximately optimize these two conditions (we fix $T=300/\Delta E$), although this still leads to a small degree of dispersion.
Figure 4: Realizations of the extended Rice-Mele model. a: In the extended Rice-Mele model, we consider in general non-zero hopping parameters between all sites. Here, an example for a system with $N=8$ sites. b: In the Rydberg system, the two sublattices are separated by a perpendicular distance $h$ and a horizontal offset $b(t)$, which can be changed in time to realize the photon pumping. Here, the nearest neighbor hoppings dominate and we exploit the dependence of the couplings on dipole moment orientation (right panel) to make the intra-sublattice hoppings ($J_{2p}$) negligible. c: Similarly, on a dense chain of atoms excited to low-lying electronic states, $J_{2}$ is put to zero by fixing a value of $a$ that depends on the angle $\theta_{\mathrm{d}}$ of the dipoles, while the rest of the hoppings (determined by $V_{ij}$) are non-zero in general. d: In a lattice of atoms coupled to a waveguide (here a cylindrical nanofiber), the atoms of sublattice $A$ are fixed forming a helix around the fiber, while the atoms of sublattice $B$ change their relative angle to sublattice $A$, $\phi^{\prime}_{1}(t)$, time-dependently to realize the pumping. Sublattice symmetry is preserved by choosing $a=\pi/\beta$, where $\beta$ is the wave vector of the light propagating inside the fiber.
@@ 4.1 Short-range hopping: Rydberg atoms
Figure 5: Photon pumping on a Rydberg chain. a: Dependence of the extended hopping parameters $\bar{J}^{\prime}$ and $\bar{J}$ with the ratio $b/a$ (all parameters are normalized to the maximal value $\bar{J}_{\text{max}}\equiv|\text{max}\{\bar{J}(t)\}|$ during the cycle). The gray shaded area represents $\delta\bar{J}<0$, where the winding number of the extended SSH model is $\nu=1$ (TP: topological phase). b: Variation of the intracell distance $b(t)$ during the pumping cycle. Importantly, twice during the cycle the extended hopping parameters coincide, such that $\delta\bar{J}=0$. c: We vary the energy offset as $\Delta(t)=\Delta_{\mathrm{max}}\sin(2\pi t/T+\pi/2)$ such that the path in parameter space encircles the topological degeneracy point ensuring the displacement of the center of mass after a completed cycle. d: The left column shows the time-integrated Berry curvature of the lower band ${\cal W}_{k}^{-}$ and quasi-momentum distribution density $|f(k)|^{2}$ (in arbitrary units). The right column shows the spatial probability density of the wave packet in real space (on the $B$ sublattice), initially (dashed line) and after its numerical evolution for ten pumping cycles (solid line). The width of the Gaussian wave packet in momentum space, $w_{k}$, is doubled from the upper ($w_{k}a=2\pi/100$) to the lower row ($w_{k}a=2\pi/50$), leaving the displacement of the photon’s center of mass $\Delta x=10\cdot 0.966a$ unchanged, but leading to an increase of the dispersion, such that the fidelity per cycle ${\cal F}(T)=0.999$ and after ten cycles ${\cal F}(10T)=0.988$ decreases to 0.997 and 0.964, respectively. Here, $T=168/\bar{J}_{\text{max}}$.
The first system we consider is an ensemble of Rydberg atoms trapped in two lattices with lattice constant $a$ and offset by a distance $b$ with respect to each other as shown in Fig. 4[42, 68]. Atoms excited to a Rydberg state are known for possessing exaggerated properties, such as long lifetimes and, most importantly, extremely large transition dipole moments between neighboring Rydberg states which can lead to strong long-range dipole-dipole interactions [69, 70]. The dynamics of the system is determined by the Hamiltonian (8), where both ground and excited states of the two-level system are Rydberg states. The dipole-dipole interaction is here given generally by the expression
@@
where $\mathbf{r}_{ij}=r_{ij}\hat{\mathbf{r}}_{ij}$ is the distance between the atoms, $\mathbf{d}=d\hat{\mathbf{d}}$ is the transition dipole moment, and $\cos{\theta_{ij}}=\hat{\mathbf{r}}_{ij}\cdot\hat{\mathbf{d}}$. To avoid breaking sublattice symmetry, we set the intra-sublattice parameters $J_{2p}$ to zero by aligning the dipoles such that they form an angle $\theta_{m}=\arccos(1/\sqrt{3})$ with the unit separation vector between atoms of the same sublattice (see Fig. 4b) [42]. Moreover, the dipole-dipole interactions (excitation hopping parameters) between neighboring atoms (lattice sites) dominate the dynamics, such that $\bar{J}^{\prime}\approx J_{1}^{\prime}$ and $\bar{J}\approx J_{1}$ with
@@
Here, $r_{1}^{\prime}=\sqrt{b^{2}+h^{2}}$, $r_{1}=\sqrt{(a-b)^{2}+h^{2}}$, $\cos\theta_{1}^{\prime}=\left(b\cos\theta_{m}-h\sin\theta_{m}\right)/r_{1}^{\prime}$ and $\cos\theta_{1}=\left[(a-b)\cos\theta_{m}+h\sin\theta_{m}\right]/r_{1}$. Adding an on-site potential offset between the even and odd sites, the Hamiltonian is very close to the standard Rice-Mele model (4). Note, however, that this is only an approximation, and the actual long-ranged character of the interactions in the Rydberg system is fully taken into account in the following in our numerical calculations.
@@ 4.1.1 Photon pumping on a Rydberg chain
Figure 6: Photon pumping on a two-level atom chain. a: Dependence of the extended hopping parameters $\bar{J}^{\prime}$ and $\bar{J}$ with the ratio $b/a$ (normalized to $\bar{J}_{\text{max}}\equiv|\text{max}\{\bar{J}(t)\}|$ during the cycle). The gray shaded area represents $\delta\bar{J}<0$, where the winding number of the extended SSH model is $\nu=1$ (TP: topological phase). b: Variation of the intracell distance $b(t)$ during the pumping cycle. c: We vary the energy offset as $\Delta(t)=\Delta_{\mathrm{max}}\sin(2\pi t/T+\pi/2)$ such that the path in parameter space encircles the topological degeneracy point ensuring the displacement of the center of mass after a completed cycle. d: The left column shows the time-integrated Berry curvature of the lower band ${\cal W}_{k}^{-}$ and quasi-momentum distribution density $|f(k)|^{2}$ (in arbitrary units). The right column shows the spatial probability density of the wave packet in real space (on the $B$ sublattice), initially (dashed line) and after its numerical evolution for ten pumping cycles (solid line). The width of the Gaussian wave packet in momentum space $w_{k}$ (the same as in Fig. 5) is doubled from the upper to the lower row, leaving the displacement of the photon’s center of mass $\Delta x=10\cdot 0.803a$ unchanged, but leading to an increase of the dispersion, such that the fidelity per cycle ${\cal F}(T)=0.999$ and after ten cycles ${\cal F}(10T)=0.986$ decreases to 0.986 and 0.890, respectively. Here, $T=190/\bar{J}_{\text{max}}$.
We simulate the dynamics starting from a wave packet on the lower band, with a Gaussian quasi-momentum distribution $|f(k)|^{2}$ of width $w_{k}$. Here, we are able to find an optimized path in parameter space that leads to a particularly flat time-integrated Berry curvature across all the FBZ. As one can also see in Figure 5d, the cyclic pumping indeed leaves the shape of the photon mostly unchanged except for its average position, which is displaced by an amount $\Delta x$. This displacement is in agreement with the one based on the time-integrated Berry curvature and the quasi-momentum distribution predicted by Equation (2). Notably, here we are able to achieve a fidelity per cycle well above $99.9\%$, which goes down as we increase the width of the wave packet in $k$-space and dispersion plays a more prominent role.
@@ 4.2 Long-range hopping: Atoms in free space
where $\gamma=\frac{d^{2}k_{0}^{3}}{3\pi\epsilon_{0}\hbar}$ is the single atom spontaneous decay rate, and $k_{0}=\omega_{0}/c=2\pi/\lambda_{0}$ is the wave number of the transition, respectively [73]. This interaction can no longer be realistically approximated to be nearest neighbor only, as it was the case in the Rydberg system. Thus, the corresponding extended Rice-Mele model (LABEL:ExtendedRM) contains all hopping parameters $J^{\prime}_{2p-1}$, $J_{2p-1}$ and $J_{2p}$. While the sublattice symmetry is in general broken in this system, we minimize this effect by choosing the distance $a$ between neighboring atoms of the same sublattice at a value $a(\theta_{d})$, where $\theta_{d}$ is the angle between the atomic dipoles and the chain, i.e. $\cos{\theta_{d}}=\hat{\mathbf{r}}_{ij}\cdot\hat{\mathbf{d}}$, such that $J_{2}=0$ (see Fig. 4c for the $\theta_{d}=\pi/2$ case). Given that the hopping parameters slowly decrease with the distance between atoms, the values of the subsequent symmetry-breaking hopping parameters $J_{4}$, $J_{6}$, are almost negligible, and the sublattice symmetry is only slightly broken.